OpenStudy (anonymous):

]Choose the equation of the horizontal line that passes through the point (-8, -7)

OpenStudy (anonymous):

for any horizontal line, always y=constant, so y=-7

jimthompson5910 (jim_thompson5910):

draw a horizontal line through (-8, -7) you'll notice that any point on this line will have a y coordinate of -7 so that's another way to see that the answer is y = -7

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

start with y = mx + b then plug in the slope to get y = (1/2)x + b after that, plug in the given point to get -4 = (1/2)*(2) + b now solve for b

OpenStudy (anonymous):

i dont know how @jim_thompson5910

jimthompson5910 (jim_thompson5910):

-4 = (1/2)*(2) + b -4 = (1/2)*(2/1) + b -4 = (1*2)/(2*1) + b -4 = 2/2 + b -4 = 1 + b ... ... ... b = ???

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

keep in mind you're subtracting 1 from both sides

jimthompson5910 (jim_thompson5910):

and -4-1 is NOT -3

OpenStudy (anonymous):

im lost @jim_thompson5910

jimthompson5910 (jim_thompson5910):

think about it like this -4 means you're 4 dollars in debt then you spend one more dollar how many dollars in debt are you now?

OpenStudy (anonymous):

-5

jimthompson5910 (jim_thompson5910):

so -4-1 = -5

jimthompson5910 (jim_thompson5910):

which means b = -5

jimthompson5910 (jim_thompson5910):

m = 1/2 b = -5

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2 ?

jimthompson5910 (jim_thompson5910):

\[\large y = mx+b\] \[\large y = \frac{1}{2}x - 5\]

jimthompson5910 (jim_thompson5910):

same thing, now m = 3 and b = -2

OpenStudy (anonymous):

last one

jimthompson5910 (jim_thompson5910):

m = slope b = y-intercept

OpenStudy (anonymous):

now what ?

jimthompson5910 (jim_thompson5910):

slope of 3 and a y-intercept of -2 ---> m = 3 and b = -2

jimthompson5910 (jim_thompson5910):

I'm sure you see what to do from here

OpenStudy (anonymous):

-2=3x+-2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you leave y as it is

jimthompson5910 (jim_thompson5910):

so it's just y = 3x - 2

OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x + 7y = -21 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

hmm you said it was the last one

jimthompson5910 (jim_thompson5910):

anyways, plug in x = 0 and solve for y to get the y-intercept plug in y = 0 and solve for x to get the x-intercept

OpenStudy (anonymous):

how ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

when you plug in x = 0, you're basically replacing x with 0

jimthompson5910 (jim_thompson5910):

then you solve for y

OpenStudy (anonymous):

idk how

jimthompson5910 (jim_thompson5910):

you've never solved equations?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

I recommend checking out this page http://www.purplemath.com/modules/solvelin.htm

jimthompson5910 (jim_thompson5910):

keep in mind that there are multiple pages and try out the examples as well

OpenStudy (anonymous):

i will but can you please just give me this last answer ? i need a good grade on this assignment

jimthompson5910 (jim_thompson5910):

sorry can't just hand out answers (it's against the rules on this site)

jimthompson5910 (jim_thompson5910):

have a look at that page

OpenStudy (anonymous):

:/ explain it to me then ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

is the page not helping at all?

jimthompson5910 (jim_thompson5910):

it should explain how to solve linear equations